Honeybees actively foraging in the field have been found to be electrically charged, due largely to air resistance as they fly. The charges carried by a bee, which can be as great as 93 pC, is thought to play a significant role in pollination -- the bee can detach grains of pollen from a distance, like a charged comb attracting bits of paper. Given the force required to detach pollen from an avocado stigma is 4.5 ✕ 10-10 N, find the maximum distance at which the electrostatic force between a bee and a grain of pollen is sufficient to detach the pollen. Treat the bee and pollen as point charges, and assume that the pollen has a charge opposite in sign and equal in magnitude to the bee.
Answer #1

Electrostatic force is given by,

Fe = k*q1*q2/r^2

here, k = 9*10^9

q1 & q2 = charges = 93 pC = 93*10^-12 C

r = distance between bee & pollen = ??

given Fe = 4.5*10^-10 N

So,

r^2 = k*q1*q2/Fe

r = sqrt(k*q1*q2/Fe)

r = sqrt[(9*10^9)*(93*10^-12)*(93*10^-12)/(4.5*10^-10)]

r = 0.4159 m

r = 41.59 cm

therefore, the maximum distance = r =42 cm

Please upvote.